# Imagined Autocurricula

### v1

@@ 1 Introduction
We believe world models can be the solution to this problem–we first train a diffusion world model from the offline samples and then use it to generate new, imagined worlds. Crucially, naively generating data from world models can lead to ineffective training data, making agent learning inefficient. This inefficiency stems from both the quality of the worlds themselves and the challenge of setting appropriate hyperparameters like imagined episode length across diverse tasks. To address this challenge, we leverage the Unsupervised Environment Design paradigm (UED) (dennis2020emergent, ))—where a teacher proposes levels that maximize student regret. Rather than training on fixed or random sets of imagined worlds, we create an Imagined Autocurricula, or iMac.
Our approach iMac uses Prioritized Level Replay (PLR, (jiang2021prioritized, )) as a UED algorithm, which we show provides a natural complement to the learned world model–the world model generates diverse potential training trajectories or "imagined environments," while PLR strategically selects subsequent training tasks from these imagined rollouts. Figure 1 illustrates the overall architecture of our approach. As we show, this prioritization process naturally induces an automatic curriculum over the generated, imagined worlds, meaning the agent is exposed to increasingly challenging training tasks. As a testbed for this paradigm, we leverage five Procgen environments with high-dimensional visual inputs, where both model-free (prudencio2023offlineRLsurvey, ) and model-based approaches (he2023surveyofflinemodelbasedreinforcement, ) struggle to transfer knowledge to unseen scenarios (mediratta2024generalization, ; lu2023challenges, ). The resulting agents have significantly stronger generalization performance than state-of-the-art baselines, demonstrating strong transfer performance on held-out environments having trained only inside a world model learned from a narrower dataset. We believe this opens the path to utilizing larger-scale, foundation world models for generally capable agents.
@@ 4.1 Procgen Benchmark
We created a mixed offline dataset capturing diverse interaction patterns to represent the type of diverse data we may see in Internet videos. Similar to “mixed” datasets in existing benchmarks (fu2021drl, ; lu2023challenges, ) it consists of three complementary components: Expert trajectories ($D_{\text{expert}}$) from fully trained PPO agents, Medium-quality trajectories ($D_{\text{medium}}$) from partially trained agents reaching approximately 50% of expert performance, and Random exploration trajectories ($D_{\text{random}}$) to ensure broad state space coverage. This mixture forms a dataset $D=D_{\text{random}}\cup D_{\text{expert}}\cup D_{\text{medium}}$ totaling 1 million transitions per game across 200 training levels. We focus on five environments from Procgen: CoinRun, Ninja, Jumper, Maze, and CaveFlyer, selected based on their poor performance in previous offline RL benchmarks (mediratta2024generalization, ) and their sparse reward structure. For each environment, we use 200 levels for training our world model while evaluating generalization on unseen test levels (201-$\infty$).
@@ 4.2 Procgen Benchmark Results
Method CoinRun Ninja Jumper Maze CaveFlyer
BC $5.31\pm 0.31$ $4.22\pm 0.21$ $3.92\pm 0.42$ $4.03\pm 0.29$ $3.12\pm 0.17$
CQL $5.12\pm 0.22$ $3.29\pm 0.32$ $2.24\pm 0.31$ $1.82\pm 0.12$ $1.89\pm 0.27$
IQL $4.96\pm 0.20$ $3.11\pm 0.31$ $2.35\pm 0.38$ $1.98\pm 0.17$ $1.45\pm 0.47$
BCQ $4.33\pm 0.14$ $3.14\pm 0.26$ $2.29\pm 0.18$ $1.83\pm 0.14$ $2.21\pm 0.61$
BCT $5.11\pm 0.10$ $4.33\pm 0.28$ $3.72\pm 0.37$ $3.98\pm 0.19$ $2.76\pm 0.22$
DT $5.03\pm 0.18$ $4.03\pm 0.31$ $3.84\pm 0.24$ $4.32\pm 0.11$ $3.04\pm 0.32$
iMac $\bm{6.20}\pm 0.14$ $\bm{5.15}\pm 0.12$ $\bm{5.78}\pm 0.21$ $\bm{5.41}\pm 0.19$ $\bm{3.87}\pm 0.22$
@@
Method CoinRun Ninja Jumper Maze CaveFlyer
WM Fixed $4.76\pm 0.19$ $3.78\pm 0.21$ $3.71\pm 0.19$ $3.92\pm 0.32$ $2.94\pm 0.13$
WM Random $5.36\pm 0.15$ $4.33\pm 0.18$ $4.11\pm 0.11$ $4.62\pm 0.15$ $3.02\pm 0.19$
iMac $\bm{6.20}\pm 0.14$ $\bm{5.15}\pm 0.12$ $\bm{5.78}\pm 0.21$ $\bm{5.41}\pm 0.19$ $\bm{3.87}\pm 0.22$
The results in Table 1 demonstrate that iMac method consistently outperforms state-of-the-art offline RL algorithms across all evaluated Procgen environments. Our approach achieves improvements of 17% on CoinRun, 19% on Ninja, 48% on Jumper, 35% on Maze, and 24% on CaveFlyer compared to the best-performing model-free baseline for each environment.
@@
Figure 2: Mean test performance across five Procgen environments. Left: Comparison of model-free baselines against our proposed iMac method. Right: World model ablation study comparing WM variants (Fixed and Random) against the full iMac approach
@@
Figure 4: Left: Difference in imagined episode length between PLR and Random baselines across five Procgen environments. Right: Example frames from early $t_{n}$ and late $t_{N}$ training stages, with red boxes representing agents and orange boxes representing rewards.
@@ 5 Related Work
World Models have evolved significantly since the first introduction of the concept of reinforcement learning within an imagined neural network environment (ha2018world, ). Early applications to Atari games through SimPLe (kaiser2024modelbasedreinforcementlearningatari, ) established the Atari 100k benchmark, highlighting the sample efficiency benefits of model-based approaches. The field advanced substantially with DreamerV2 (hafner2020mastering, ), which pioneered learning directly from the latent space of a recurrent state space model (RSSM), while DreamerV3 (hafner2024masteringdiversedomainsworld, ) achieved human-level performance across diverse domains using consistent hyperparameters. Recent architectural innovations include transformer-based approaches such as TWM (robine2023smaller, ) and STORM (zhang2020exemplarnormalizationlearningdeep, ), which adapt DreamerV2’s and DreamerV3’s RSSM frameworks respectively, employing different tokenization strategies. IRIS (micheli2023transformerssampleefficientworldmodels, ) represents an alternative approach, constructing a language of image tokens through a discrete autoencoder and composing these tokens temporally using an autoregressive transformer. Most recently, DIAMOND (alonso2024diffusionworldmodelingvisual, ) introduced diffusion models as an alternative to discrete latent representations for world modeling, demonstrating that improved visual fidelity can lead to better agent performance, particularly for tasks where fine visual details are critical. There has also been work using world models to generalize to unseen tasks (ball21a, ; lee2020context, ), however these were focused on a domain randomization setting (e.g. our “random” baseline) rather than an automated curriculum. By contrast, our work builds upon these foundations and leverages PLR to generate an emergent curriculum, which subsequently improves generalization to unseen tasks.
@@
Curriculum Learning in RL applies the principle of gradually increasing task difficulty to reinforcement learning, improving both training efficiency and generalization capabilities (Bengio2009CurriculumL, ; matiisen2017teacherstudentcurriculumlearning, ). Determining learning priorities is central to effective curriculum design, with various measures including reward signals (andrychowicz2018hindsightexperiencereplay, ), learning progress (portelas20a, ), TD-errors (schaul2016prioritizedexperiencereplay, ), and state novelty (savinov2019episodiccuriosityreachability, ). Recent advances in automated curriculum learning (graves2017automatedcurriculumlearningneural, ) have reduced reliance on manual design, with methods like Prioritized Level Replay (PLR) (jiang2021prioritized, ) demonstrating that targeting environments with the highest learning potential creates an emergent curriculum that progressively challenges the agent at the frontier of its capabilities. Our work applies this curriculum-based approach to world models, creating an imagined autocurriculum that operates entirely within generated rollouts from a fixed world model trained on offline data, enabling effective generalization without requiring explicit environment control or online interaction.
@@ 7 Conclusion
We introduced iMac, a novel approach combining diffusion-based world models with automatic curriculum learning to train general agents. Our method leverages offline data to generate diverse imagined environments and employs Prioritized Level Replay to create an emergent curriculum that adapts to the agent’s capabilities. Experimental results across five Procgen environments demonstrate that iMac consistently outperforms state-of-the-art offline RL algorithms and world model baselines, achieving improvements of up to 48% compared to model-free approaches and 56% over fixed-horizon world models. The emergent curriculum proves particularly effective for environments requiring extended planning horizons and complex navigation. Our work demonstrates that strong transfer performance on held-out environments is possible using only imagined trajectories from a world model trained on a limited dataset. This finding opens a promising path toward utilizing larger-scale foundation world models for developing generally capable agents that can handle novel task variations.

### v2

@@ 1 Introduction
We believe world models can be the solution to this problem–we first train a diffusion world model from the offline samples and then use it to generate new, imagined worlds. Crucially, naively generating data from world models can lead to ineffective training data, making agent learning inefficient. This inefficiency stems from both the quality of the worlds themselves and the challenge of setting appropriate hyperparameters like imagined episode length across diverse tasks. To address this challenge, we leverage the Unsupervised Environment Design paradigm (UED) (dennis2020emergent, ))—where a teacher proposes levels that maximize student regret. Rather than training on fixed or random sets of imagined worlds, we create an Imagined Autocurricula, or iMac. Importantly, our framework is world-model agnostic—while we use a diffusion world model as our foundation, our autocurriculum approach can be applied to any world model architecture that provides reward and next-token predictions. Our contribution lies not in advancing world model architectures, but in demonstrating how UED principles can effectively guide agent training within learned world models from offline mixed dataset.
Our approach iMac uses Prioritized Level Replay (PLR, (jiang2021prioritized, )) as a UED algorithm, which we show provides a natural complement to the learned world model–the world model generates diverse potential training trajectories or "imagined environments," while PLR strategically selects subsequent training tasks from these imagined rollouts. Figure 1 illustrates the overall architecture of our approach. As we show, this prioritization process naturally induces an automatic curriculum over the generated, imagined worlds, meaning the agent is exposed to increasingly challenging training tasks. As a testbed for this paradigm, we leverage seven Procgen environments with high-dimensional visual inputs, where both model-free (prudencio2023offlineRLsurvey, ) and model-based approaches (he2023surveyofflinemodelbasedreinforcement, ) struggle to transfer knowledge to unseen scenarios (mediratta2024generalization, ; lu2023challenges, ). The resulting agents have significantly stronger generalization performance than state-of-the-art baselines, demonstrating strong transfer performance on held-out environments having trained only inside a world model learned from a narrower dataset. We believe this opens the path to utilizing larger-scale, foundation world models for generally capable agents.
@@ 4.1 Procgen Benchmark
We created a mixed offline dataset capturing diverse interaction patterns to represent the type of diverse data we may see in Internet videos. Similar to “mixed” datasets in existing benchmarks (fu2021drl, ; lu2023challenges, ) it consists of three complementary components: Expert trajectories ($D_{\text{expert}}$) from fully trained PPO agents, Medium-quality trajectories ($D_{\text{medium}}$) from partially trained agents reaching approximately 50% of expert performance, and Random exploration trajectories ($D_{\text{random}}$) to ensure broad state space coverage. This mixture forms a dataset $D=D_{\text{random}}\cup D_{\text{expert}}\cup D_{\text{medium}}$ totaling 1 million transitions per game across 200 training levels. We focus on seven environments from Procgen: CoinRun, Ninja, Jumper, Maze, and CaveFlyer, selected based on their poor performance in previous offline RL benchmarks (mediratta2024generalization, ) and their sparse reward structure. For each environment, we use 200 levels for training our world model while evaluating generalization on unseen test levels (201-$\infty$).
@@ 4.2 Procgen Benchmark Results
Method CoinRun Ninja Jumper Maze CaveFlyer Heist Miner
BC $5.31\pm 0.31$ $4.22\pm 0.21$ $3.92\pm 0.42$ $4.03\pm 0.29$ $3.12\pm 0.17$ $2.1\pm 0.22$ $4.68\pm 0.11$
CQL $5.12\pm 0.22$ $3.29\pm 0.32$ $2.24\pm 0.31$ $1.82\pm 0.12$ $1.89\pm 0.27$ $0.3\pm 0.09$ $0.32\pm 0.09$
IQL $4.96\pm 0.20$ $3.11\pm 0.31$ $2.35\pm 0.38$ $1.98\pm 0.17$ $1.45\pm 0.47$ $0.27\pm 0.04$ $0.5\pm 0.08$
BCQ $4.33\pm 0.14$ $3.14\pm 0.26$ $2.29\pm 0.18$ $1.83\pm 0.14$ $2.21\pm 0.61$ $0.44\pm 0.12$ $0.42\pm 0.13$
BCT $5.11\pm 0.10$ $4.33\pm 0.28$ $3.72\pm 0.37$ $3.98\pm 0.19$ $2.76\pm 0.22$ $1.5\pm 0.23$ $1.08\pm 0.10$
DT $5.03\pm 0.18$ $4.03\pm 0.31$ $3.84\pm 0.24$ $4.32\pm 0.11$ $3.04\pm 0.32$ $82\pm 0.09 1.1$$11\pm 0.09$
iMac $\bm{6.20}\pm 0.14$ $\bm{5.15}\pm 0.12$ $\bm{5.78}\pm 0.21$ $\bm{5.41}\pm 0.19$ $\bm{3.87}\pm 0.22$ $\bm{2.91}\pm 0.15$ $\bm{5.61}\pm 0.13$
@@
Method CoinRun Ninja Jumper Maze CaveFlyer Heist Miner
WM Fixed $4.76\pm 0.19$ $3.78\pm 0.21$ $3.71\pm 0.19$ $3.92\pm 0.32$ $2.94\pm 0.13$ $2.05\pm 0.14$ $3.18\pm 0.10$
WM Random $5.36\pm 0.15$ $4.33\pm 0.18$ $4.11\pm 0.11$ $4.62\pm 0.15$ $3.02\pm 0.19$ $2.19\pm 0.18$ $4.55\pm 0.15$
iMac $\bm{6.20}\pm 0.14$ $\bm{5.15}\pm 0.12$ $\bm{5.78}\pm 0.21$ $\bm{5.41}\pm 0.19$ $\bm{3.87}\pm 0.22$ $\bm{2.91}\pm 0.15$ $\bm{5.61}\pm 0.13$
The results in Table 1 demonstrate that iMac method consistently outperforms state-of-the-art offline RL algorithms across all evaluated Procgen environments. Our approach achieves improvements of 17% on CoinRun, 19% on Ninja, 48% on Jumper, 35% on Maze, 24% on CaveFlyer, 38% on Heist, and 19% on Miner compared to the best-performing model-free baseline for each environment.
@@
Figure 2: Mean test performance across seven Procgen environments. Left: Comparison of model-free baselines against our proposed iMac method. Right: World model ablation study comparing WM variants (Fixed and Random) against the full iMac approach
@@
Figure 4: Left: Difference in imagined episode length between PLR and Random baselines across seven Procgen environments. Right: Example frames from early $t_{n}$ and late $t_{N}$ training stages, with red boxes representing agents and orange boxes representing rewards.
@@ 5 Related Work
World Models have evolved significantly since the first introduction of the concept of reinforcement learning within an imagined neural network environment (ha2018world, ). Early applications to Atari games through SimPLe (kaiser2024modelbasedreinforcementlearningatari, ) established the Atari 100k benchmark, highlighting the sample efficiency benefits of model-based approaches. The field advanced substantially with DreamerV2 (hafner2020mastering, ), which pioneered learning directly from the latent space of a recurrent state space model (RSSM), while DreamerV3 (hafner2024masteringdiversedomainsworld, ) achieved human-level performance across diverse domains using consistent hyperparameters. Recent architectural innovations include transformer-based approaches such as TWM (robine2023smaller, ) and STORM (zhang2020exemplarnormalizationlearningdeep, ), which adapt DreamerV2’s and DreamerV3’s RSSM frameworks respectively, employing different tokenization strategies. IRIS (micheli2023transformerssampleefficientworldmodels, ) represents an alternative approach, constructing a language of image tokens through a discrete autoencoder and composing these tokens temporally using an autoregressive transformer. Most recently, DIAMOND (alonso2024diffusionworldmodelingvisual, ) introduced diffusion models as an alternative to discrete latent representations for world modeling, demonstrating that improved visual fidelity can lead to better agent performance, particularly for tasks where fine visual details are critical. Recent advances in video generation, including 3D DiT architectures like CogVideoX (yang2025cogvideoxtexttovideodiffusionmodels, ) and Cosmos (nvidia2025cosmosworldfoundationmodel, ), offer promising alternatives for world modeling. While we employ a 2D UNet for computational efficiency, our curriculum learning framework is architecture- agnostic and could leverage these more advanced architectures as they become computationally more feasible. There has also been work using world models to generalize to unseen tasks (ball21a, ; lee2020context, ), however these were focused on a domain randomization setting (e.g. our “random” baseline) rather than an automated curriculum. By contrast, our work builds upon these foundations and leverages PLR to generate an emergent curriculum, which subsequently improves generalization to unseen tasks.
@@
Curriculum Learning in RL applies the principle of gradually increasing task difficulty to reinforcement learning, improving both training efficiency and generalization capabilities (Bengio2009CurriculumL, ; matiisen2017teacherstudentcurriculumlearning, ). Determining learning priorities is central to effective curriculum design, with various measures including reward signals (andrychowicz2018hindsightexperiencereplay, ), learning progress (portelas20a, ), TD-errors (schaul2016prioritizedexperiencereplay, ), and state novelty (savinov2019episodiccuriosityreachability, ). Recent advances in automated curriculum learning (graves2017automatedcurriculumlearningneural, ) have reduced reliance on manual design, with methods like Prioritized Level Replay (PLR) (jiang2021prioritized, ) demonstrating that targeting environments with the highest learning potential creates an emergent curriculum that progressively challenges the agent at the frontier of its capabilities. While we employ PLR in our implementation, our framework is compatible with other UED algorithms including ACCEL (parkerholder2023evolvingcurricularegretbasedenvironment, ), which uses evolution strategies for curriculum design; ADD (chung2024adversarialenvironmentdesignregretguided, ), which leverages diffusion models for adversarial environment generation. These methods could potentially be integrated into our framework, highlighting the modularity of our approach. Our work applies this curriculum-based approach to world models, creating an imagined autocurriculum that operates entirely within generated rollouts from a fixed world model trained on offline data, enabling effective generalization without requiring explicit environment control or online interaction.
@@ 7 Conclusion
We introduced iMac, a novel approach combining diffusion-based world models with automatic curriculum learning to train general agents. Our method leverages offline data to generate diverse imagined environments and employs Prioritized Level Replay to create an emergent curriculum that adapts to the agent’s capabilities. Experimental results across seven Procgen environments demonstrate that iMac consistently outperforms state-of-the-art offline RL algorithms and world model baselines, achieving improvements of up to 48% compared to model-free approaches and 56% over fixed-horizon world models. The emergent curriculum proves particularly effective for environments requiring extended planning horizons and complex navigation. Our work demonstrates that strong transfer performance on held-out environments is possible using only imagined trajectories from a world model trained on a limited dataset. This finding opens a promising path toward utilizing larger-scale foundation world models for developing generally capable agents that can handle novel task variations.
@@
Appendix A Model Architectures
Diffusion World Model
The diffusion model $D_{\theta}$ employs a standard 2D U-Net architecture [30] for next-frame prediction. The model conditions on a history window of 4 previous frames and their corresponding actions, along with the diffusion timestep $\tau$. We implement observation conditioning through frame stacking along the channel dimension, while action and diffusion time conditioning are incorporated via adaptive group normalization layers [31] within the U-Net’s residual blocks.
Reward and Termination Predictor
The reward/termination model $R_{\psi}$ utilizes a shared backbone architecture with separate prediction heads for reward and termination signals. Crucially, we employ an ensemble of 10 prediction heads to enable uncertainty quantification during imagination rollouts. The model processes sequences of frames and actions through convolutional residual blocks [32] followed by an LSTM cell [35, 34]. During inference, the ensemble predictions are aggregated to compute both mean estimates and uncertainty measures, providing robustness against prediction errors in long-horizon imagined trajectories. Prior to imagination, we perform burn-in [74] using the conditioning frames and actions to properly initialize the LSTM’s hidden and cell states.
Actor-Critic Network
The policy $\pi_{\phi}$ and value network $V_{\phi}$ share a common feature extraction backbone with separate output layers. We refer to this combined architecture as the actor-critic network $(\pi,V)_{\phi}$, though technically $V$ represents a state-value function rather than an action-value critic. The network processes individual frames through a convolutional encoder consisting of four residual blocks with $\times2$ max-pooling (stride 2) between blocks. Each residual block contains a main path with group normalization [75], SiLU activation [76], and $\times3$ convolution (stride 1, padding 1). The convolutional features are then processed by an LSTM cell to maintain temporal context. Similar to the reward model, we initialize the LSTM states through burn-in on conditioning frames before beginning the imagination procedure.
Please refer to Table 3 below for hyperparameter values.
Table 3: Architecture Parameters for iMAC.
Hyperparameter Value
Diffusion Model ($D_{\theta}$)
Observation conditioning mechanism Frame stacking
Action conditioning mechanism Adaptive Group Normalization
Diffusion time conditioning mechanism Adaptive Group Normalization
Residual blocks layers [2, 2, 2, 2]
Residual blocks channels [64, 64, 64, 64]
Residual blocks conditioning dimension 256
Reward/Termination Model ($R_{\psi}$)
Action conditioning mechanism Adaptive Group Normalization
Residual blocks layers [2, 2, 2, 2]
Residual blocks channels [32, 32, 32, 32]
Residual blocks conditioning dimension 128
LSTM dimension 512
Actor-Critic Model ($\pi_{\phi}$ and $V_{\phi}$)
Residual blocks layers [1, 1, 1, 1]
Residual blocks channels [32, 32, 64, 64]
LSTM dimension 512
Appendix B World Model Training Parameters
Table 4: Hyperparameters for iMAC.
Hyperparameter Value
Training loop
Number of epochs 1000
Training steps per epoch 100
Batch size 32
Environment steps per epoch 100
Epsilon (greedy) for collection 0.01
RL hyperparameters
Imagination horizon ($H$) Fixed:15, Random:[5-22]
Discount factor ($\gamma$) 0.985
Entropy weight ($\eta$) 0.001
$\lambda$-returns coefficient ($\lambda$) 0.95
Sequence construction during training
For $D_{\theta}$, number of conditioning observations and actions ($L$) 4
For $R_{\psi}$, burn-in length ($B_{R}$), set to $L$ in practice 4
For $R_{\psi}$, training sequence length ($B_{R}+H$) Fixed:19, Random[9-26]
For $\pi_{\phi}$ and $V_{\phi}$, burn-in length ($B_{\pi,V}$), set to $L$ in practice 4
Optimization
Optimizer AdamW
Learning rate 4e-5
Epsilon 1e-8
Weight decay ($D_{\theta}$) 1e-2
Weight decay ($R_{\psi}$) 1e-2
Weight decay ($\pi_{\phi}$ and $V_{\phi}$) 5e-5
Diffusion Sampling
Method Euler
Number of steps 5
Environment
Image observation dimensions $\times64$$\times$3
Action space Discrete (up to 18 actions)
Frameskip 4
Max noop 30
Termination on life loss True
Reward clipping $\{0,1\}$
PLR
Stalaness Factor 0.1
Buffer Size 2500
Appendix C Baseline World Model and Actor-Critic Training Algorithm
Procedure training_loop($D$)
 // $D$ is a fixed offline dataset;
 Phase 1: Train Diffusion Model;
 for epochs_diffusion do do
    for steps_diffusion_model do do
         update_diffusion_model($D$);
      end for
   end for
 // Freeze Diffusion Model $D_{\theta}$;
 Phase 2: Train Reward/Termination Model;
 for epochs_reward do do
    for steps_reward_end_model do do
         update_reward_end_model($D$);
      end for
   end for
 // Freeze Reward/Termination Model $R_{\psi}$;
 Phase 3: Train Actor-Critic with Frozen Models;
 for epochs_actor_critic do do
    for steps_actor_critic do do
         update_actor_critic($D$);
       // Uses frozen $D_{\theta}$ and $R_{\psi}$
      end for
   end for
Procedure update_diffusion_model($D$)
   Sample sequence $(x^{0}_{t-L+1},a_{t-L+1},\ldots,x^{0}_{t},a_{t},x^{0}_{t+1})\sim D$;
   Sample $(\sigma)\sim\mathcal{N}(F_{mean},F^{2}_{std}) // log-$normal sigma distribution from EDM;
   Define $\tau:=\sigma$ // default identity schedule from EDM;
   Sample $x^{1}_{t+1}\sim\mathcal{N}(x^{0}_{t+1},\sigma^{2}I)$ // Add independent Gaussian noise;
   Compute $x^{0}_{t+1}=D_{\theta}(x^{1}_{t+1},\tau,x^{0}_{t-L+1},a_{t-L+1},\ldots,x^{0}_{t},a_{t})$;
[…]
   Update $D_{\theta}$;
Procedure update_reward_end_model($D$)
   Sample indices $I:=\{t,\ldots,t+L+H-1\}$ // burn-in + imagination horizon;
   Sample sequence $(x^{0}_{i},a_{i},r_{i},d_{i})_{i\in I}\sim D$;
   Initialize $h=c=0$ // LSTM hidden and cell states;
 for $i\in I$ do do
      Compute $\hat{r}_{i},\hat{d}_{i},h,c=R_{\psi}(x_{i},a_{i},h,c)$;
   end for
  Compute $\mathcal{L}(\psi)=\sum_{i\in I}\text{CE}(r_{i},\text{sign}(r_{i}))+\text{CE}(d_{i},\hat{d}_{i})$ // CE: cross-entropy loss;
   Update $R_{\psi}$;
Procedure update_actor_critic($D$)
   Sample initial buffer $(x^{0}_{t-L+1},a_{t-L+1},\ldots,x^{0}_{t})\sim D$;
   Burn-in buffer with $R_{\psi},\pi_{\phi}$ and $V_{\phi}$ to initialize LSTM states;
 for $i=t$ to t+H1t+H-1 do do
      Sample $a_{i}\sim\pi_{\phi}(a_{i}|x^{0}_{i})$;
      Sample reward $r_{i}$ and termination $d_{i}$ with frozen $R_{\psi}$;
      Sample next observation $x^{0}_{i+1}$ by simulating reverse diffusion process with frozen $D_{\theta}$;
   end for
  Compute $V_{\phi}(x_{i})$ for $i=t,\ldots,t+H$;
   Compute RL losses $\mathcal{L}_{V}(\phi)$ and $\mathcal{L}_{\pi}(\phi)$;
   Update $\pi_{\phi}$ and $V_{\phi}$;
Algorithm 1 iMac (Sequential Offline Training)
Procedure train_actor_critic_PLR($D$, $B$)
 // $D$ is the offline dataset, $B$ is the prioritized buffer;
 for epochs_actor_critic do do
    // Update buffer with probability 0.5;
    if random() <0.5<0.5 then then
         update_PLR_buffer($D$, $B$);
      end if
    for steps_actor_critic do do
       // Train actor-critic using only buffer data;
         update_actor_critic_from_buffer($B$);
      end for
   end for
Procedure update_PLR_buffer($D$, $B$)
 // Sample horizon length randomly;
   Sample initial sequence $(x^{0}_{t-L+1},a_{t-L+1},\ldots,x^{0}_{t})\sim D$;
   Burn-in buffer with $R_{\psi},\pi_{\phi}$ and $V_{\phi}$ to initialize LSTM states;
 // Imagine trajectory using frozen models;
   Initialize trajectory $\tau=\{\}$;
 for $i=t$ to t+H1t+H-1 do do
      Sample $a_{i}\sim\pi_{\phi}(a_{i}|x^{0}_{i})$;
      Sample reward $r_{i}$ and termination $d_{i}$ with frozen $R_{\psi}$;
      Sample next observation $x^{0}_{i+1}$ by simulating reverse diffusion process with frozen $D_{\theta}$;
      Compute TD-error $\delta_{i}=r_{i}+\gamma V_{\phi}(x^{0}_{i+1})-V_{\phi}(x^{0}_{i})$;
    $\tau\leftarrow\tau\cup\{(x^{0}_{i},a_{i},r_{i},d_{i},x^{0}_{i+1},\delta_{i})\}$;
    if $d_{i}=1$ then then
       break;
      end if
   end for
 // Calculate PLR score as positive mean TD-error;
   Compute PLR score $s_{\tau}=\text{mean}(\max(\delta_{i},0)\text{ for }\delta_{i}\text{ in }\tau)$;
 // Update buffer if score is higher than any existing entry;
 if $s_{\tau}>\min(s_{b}\text{ for }b\in B)$ then then
      Replace lowest-scoring entry in $B$ with $(\tau,s_{\tau})$;
   end if
Procedure update_actor_critic_from_buffer($B$)
 // Sample trajectory from buffer based on priorities;
   Sample trajectory $\tau\sim B$ according to PLR scores;
 // Extract states, actions, rewards, etc.;
   Extract $(x^{0}_{i},a_{i},r_{i},d_{i},x^{0}_{i+1})$ from $\tau$;
 // Compute values for all states in trajectory;
   Compute $V_{\phi}(x^{0}_{i})$ for all states in $\tau$;
 // Compute advantage estimates and returns;
   Compute advantages $A_{i}$ and returns $G_{i}$ for all transitions;
 // Update actor and critic;
   Compute RL losses $\mathcal{L}_{V}(\phi)=\sum_{i}(V_{\phi}(x^{0}_{i})-G_{i})^{2}$;
   Compute $\mathcal{L}_{\pi}(\phi)=-\sum_{i}\log\pi_{\phi}(a_{i}|x^{0}_{i})\cdot A_{i}$;
   Update $\pi_{\phi}$ and $V_{\phi}$;
Algorithm 2 iMac: Actor-Critic Training with PLR (Prioritized Level Replay)
Appendix D Dataset Collection Details
We constructed our offline dataset by collecting trajectories from 200 procedurally generated levels for each of the five Procgen environments (CoinRun, Ninja, Jumper, Maze, and CaveFlyer), totaling 1 million environment steps per game. We explicitly constrained data collection to levels 0-199 for each environment, ensuring complete isolation from the test levels (200+) used for evaluation, thereby preventing any potential data leakage and maintaining a strict train-test split that properly assesses generalization to truly unseen procedurally generated scenarios.Following and extending the benchmark protocol from the latest work [20], we created a mixed-quality dataset with equal proportions (1/3 each) of expert, medium, and random trajectories. Expert trajectories were collected from fully-trained PPO agents [35], medium trajectories from agents achieving 50% of expert performance, and random trajectories from uniformly random action selection. Expert trajectories were collected from PPO agents trained until convergence (when learning plateaued) rather than for a fixed number of steps, ensuring optimal performance for each game’s unique characteristics.This mixed composition introduces additional challenges compared to standard offline RL benchmark [20] by reducing the proportion of successful episodes. The dataset is stored in standard (s, a, r, s’) format where observations s are 64×64×3 RGB images, actions a vary by environment-specific action spaces, and rewards are sparse—appearing only at episode termination in all five selected environments. This extreme sparsity in reward signals presents a particularly challenging scenario for world model learning, as the model must learn long-horizon dynamics without intermediate reward guidance. The hyperparameters of behavior policy are provided at Appendix F.
Appendix E Analysis
E.1 Emergent Curriculum Effect
[…]
E.2 Component Contributions
The ablation study in Table 2 reveals the critical importance of combining both PLR and random horizon sampling. While fixed-horizon world models show reasonable performance (e.g., 3.71 on Jumper), incorporating random horizons improves results by 11% (4.11), and adding PLR-based prioritization yields an additional 41% improvement (5.78). This suggests that the success of iMac stems from the synergistic effect of (1) diverse training experiences through random horizons preventing overfitting to fixed-length dynamics, and (2) intelligent prioritization that focuses computational resources on the most informative imagined trajectories.
E.3 Generalization Mechanisms
The superior generalization of iMac can be attributed to its ability to generate and prioritize diverse imagined experiences beyond the offline dataset distribution. By training exclusively on imagined rollouts from a world model learned on 200 levels, our agents achieve strong performance on unseen test levels (201+), with improvements ranging from 17% to 48% over model-free baselines. This suggests that the combination of world models and autocurricula enables agents to explore a richer space of potential scenarios than what exists in the original offline data, effectively augmenting the training distribution in a principled manner.
Appendix F Baseline Algorithm Hyperparameters
We conducted comprehensive hyperparameter optimization for all baseline algorithms using grid search following the previous benchmark work [20]. All model-free offline RL methods (BCQ, CQL, IQL) and behavior cloning (BC) shared identical encoder architectures, utilizing ResNet for visual feature extraction from the 64×64×3 Procgen observations. Experiments were executed on NVIDIA RTX 4090 GPUs with 24GB memory. For behavior cloning (BC), we explored batch sizes $\in\{64,128,256,512\}$ and learning rates $\in\{5\times 10^{-3},1\times 10^{-4},5\times 10^{-4},6\times 10^{-5}\}$. The Q-learning based methods (BCQ, CQL, IQL) employed dual-network architectures with target networks, where we investigated both Polyak averaging and direct weight copying strategies. For direct copying, we tested update frequencies $\in\{1,100,1000\}$, while for Polyak averaging, we examined $\tau\in\{0.005,0.5,0.99\}$. Algorithm-specific hyperparameters included: BCQ’s action selection threshold $\in\{0.3,0.5,0.7\}$; CQL’s regularization coefficient $\alpha\in\{0.5,1.0,4.0,8.0\}$; and IQL’s temperature $\in\{3.0,7.0,10.0\}$ with expectile weights $\in\{0.7,0.8,0.9\}$. For sequence modeling approaches (DT, BCT), we additionally optimized context lengths $\in\{5,10,30,50\}$. Decision Transformer utilized return-to-go multipliers $\in\{1,5\}$, where the maximum return was determined from training data and scaled accordingly. Both DT and BCT maintained a dropout rate of 0.1 following [43]. Table  5 presents the final selected hyperparameters achieving optimal performance on our 1M-step mixed dataset.
Table 5: Final hyperparameter configurations for baseline algorithms on 1M mixed dataset
Algorithm Hyperparameter Value
BC Learning Rate 0.0005
Batch Size 256
BCT Learning Rate 0.0005
Batch Size 512
Context Length 30
Eval Return Multiplier 0
DT Learning Rate 0.0005
Batch Size 512
Context Length 10
Eval Return Multiplier 5
BCQ Learning Rate 0.0005
Batch Size 256
Target Model Weight Update Direct copy
$\tau$ -
Target Update Frequency 1000
Threshold 0.5
CQL Learning Rate 0.0005
Batch Size 256
Target Model Weight Update Direct copy
$\tau$ -
Target Update Frequency 1000
Alpha 4.0
IQL Learning Rate 0.0005
Target Model Weight Update Direct copy
Batch Size 512
$\tau$ -
Target Update Frequency 100
Temperature 3.0
Expectile 0.8
Appendix G Training Time Profile
Table 6 presents a comprehensive breakdown of computational time requirements for training IMAC 1 game per seed.
Table 6: Detailed computational time analysis for iMac training pipeline. Measurements were obtained using an NVIDIA RTX 4090 GPU with hyperparameters detailed in previous sections. These timing results serve as representative benchmarks, with actual runtimes varying based on hardware configuration, environment setup, and specific training conditions.
Component Time (hours)
World Model Training (Offline)
   Diffusion model 6.8
   Reward/Termination model 5.4
Agent Training
   Total (1000 epochs) 75.0
   Per epoch (avg) 0.075
Total Training Time 87.2